OpenStudy (aaaa):

Arnold went shopping at the Link with $80 to spend. He spent $12 on a belt and $20 for a shirt. How much can he spend on a pair of pants? A. x < 68 B. x < 58 C. x < 48 D. x < 38

OpenStudy (anonymous):

12+20=32, so he spent $32 in total 80-32=48, so he has $48 left over to spend on pants

OpenStudy (anonymous):

both a and b are correct

OpenStudy (anonymous):

he can only spend $48.00 more because he already spent $32.00. 32 + 48 = 80

OpenStudy (anonymous):

80 - 32 = 48